# Quokka: Accelerating Program Verification with LLMs via Invariant Synthesis

### v2

@@ Abstract
Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge. We investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants. We introduce Quokka, a first-order and effective framework for LLM-based invariant synthesis that provides sound evaluation while achieving state-of-the-art speedup results. Unlike prior work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure. We construct a benchmark of 866 instances and evaluate 9 state-of-the-art LLMs across multiple model families. Our results show that Quokka consistently outperforms all prior LLM-based verifiers: achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach. We further demonstrate that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
Machine Learning, ICML
@@ 1 Introduction
Loop invariants are conditions that hold before and after each loop iteration, and they are central to deductive program verification. To accelerate program verification, loop invariants must not only be correct but also sufficiently strong to prove the assertions. Generating correct invariants is relatively easy, since any universally true condition qualifies. However, only strong invariants can reduce verification effort and lead to a speedup. For example, in Figure˜1, the invariant $x>0$ is correct but not strong enough to prove the final assertion $x\neq 700$, whereas $x\equiv 3\pmod{7}$ is both correct and sufficiently strong.
@@
A series of follow-up works have proposed LLM-based verifiers. Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs. This line of work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants. For instance, LaM4Inv (Wu et al., 2024a) uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through complex logical operations (conjunctions or disjunctions) to construct valid loop invariants, rather than using them directly. LOOPY (Kamath et al., 2024) employs a customized Houdini algorithm with iterative refinement to filter candidate invariants. LEMUR (Wu et al., 2024b) develops a backtracking and sequential algorithm. While these approaches demonstrate the potential of LLM-based verification, their complexity raises a question: can we design a simpler, principled, and more effective decision procedure for LLM-based verifiers?
@@
Figure 1: Illustration of Quokka’s evaluation pipeline. The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3). The verification procedure then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
In this work, we introduce Quokka, a framework for LLM-based invariant synthesis that accelerates program verification. Quokka provides a sound approach to evaluate LLMs’ capabilities in generating invariants while achieving state-of-the-art speedup results compared with prior LLM-based verifiers. We formalize our methodology as a proof calculus, providing a rigorous foundation for building LLM-based verifiers. Instead of checking the correctness of invariants based on string-matching, we adopt a verifier-based approach that directly queries the underlying solver. Unlike previous work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure. Our approach enables both soundness guarantees and precise measurement of performance speedup, allowing us to systematically evaluate whether LLM-generated invariants can accelerate verification in practice.
To support systematic comparison across solvers and LLMs, we construct a dataset of 866 instances derived from the most recent edition of the software verification competition SV-COMP (Beyer and Strejček, 2025). To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date, compared with datasets used in previous work. Quokka consistently outperforms all prior LLM-based verifiers. When comparing against LEMUR (the previous best LLM-based verifier with the same underlying solver), Quokka achieves speedups of at least 1.2× (i.e., verification completes at least 20% faster than the baseline solver) on 81 instances compared to LEMUR’s 39 instances. For more substantial speedups of at least 2.0× (i.e., verification completes at least twice as fast), Quokka reaches 51 instances compared to LEMUR’s 22 instances.
@@
We evaluate 9 state-of-the-art LLMs, including GPT models from OpenAI, Claude models from Anthropic, and models from the Qwen and LLaMA families. Our experiments reveal that gpt-5.2 achieves the strongest speedup results among all evaluated models. Beyond model evaluation, we investigate techniques to enhance invariant generation capabilities. We construct a training dataset of 3589 instances using a verifier-based filtering approach that ensures data quality. Our experiments demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification. Specifically, Best-of-8 sampling achieves speedups of at least 1.2× on 99 instances, representing a 22% improvement over single-sample generation, which achieves such speedups on only 81 instances.
@@
  • We introduce a decision procedure for LLM-based invariant synthesis that provides sound evaluation and achieves the best speedup results compared with prior LLM-based verifiers.
  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 state-of-the-art LLMs across multiple model families.
@@ 2 Related Work
A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan and Qadeer, 2002; Lahiri and Bryant, 2007; Hojjat and Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot and Cousot, 1977; Cousot and Halbwachs, 1978; Cousot and Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala and McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich and Bodík, 2018). Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019). Daikon executes the program, observes runtime values, and reports properties that consistently hold over the observed executions. However, such invariants may fail to generalize to all possible executions, thereby compromising soundness. Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
@@
Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014, 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020). More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a, b, c), giving rise to a series of work that explore using LLMs for finding invariants. (Pei et al., 2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation. Various techniques have been proposed to couple LLMs with symbolic solvers, including ranking LLM-generated invariants (Chakraborty et al., 2023), the “query-filter-reassemble” strategy of LaM4Inv (Wu et al., 2024a), the back-tracking algorithm in LEMUR (Wu et al., 2024b), and Loopy’s integration of the classic Houdini algorithm (Kamath et al., 2024). On the dataset side, (Liu et al., 2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach. In contrast, our work provides a simple and sound evaluation procedure for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
@@ 3.1 Preliminary
3.2 Verifier-Based Decision Procedure
We formalize our verifier-based procedure for assessing candidate invariants. Let $P$ denote a program. A property is written as $p=\langle\varphi,\ell\rangle$, where $\varphi$ is a state predicate and $\ell$ is a program location. For a finite set $A$ of properties, let $\mathrm{Asm}(P,A)$ be the program obtained from $P$ by inserting assume statements for all elements of $A$. An execution of $\mathrm{Asm}(P,A)$ that reaches a location where an assumption is violated terminates immediately. We write $P\models_{A}p$ to indicate that all executions of $\mathrm{Asm}(P,A)$ satisfy the assertion $p$. The notation $P\models p$ abbreviates $P\models_{\varnothing}p$. Since assumptions restrict behaviors, if $P\not\models_{A}p$ for some $A$, then necessarily $P\not\models p$.
@@ 3.2 Verifier-Based Decision Procedure
The verification task specifies a target property $p^{\star}=\langle\varphi^{\star},\ell^{\star}\rangle$. Given $P$ and $p^{\star}$, a large language model proposes a candidate invariant $q=\langle\psi,\ell\rangle$, typically at a loop header. To evaluate the utility of $q$, the procedure issues two verifier queries:
@@
The query $d_{a}$ verifies whether $q$ is a correct invariant, while the query $d_{b}$ determines whether the target property holds under the assumption that $q$ is true. An illustration of these two verifier queries is provided in Figure˜1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure). The outcome of the procedure is expressed as a judgment

### v3

@@ Abstract
Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge. We investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants. We introduce Quokka, an evaluation-oriented framework for LLM-based invariant synthesis that provides sound evaluation and achieves state-of-the-art performance. Unlike prior work that treats LLM outputs as noisy symbolic material requiring substantial post-processing, Quokka adopts a simpler and evaluation-centric design that directly validates whether each LLM-generated invariant helps prove the target assertion. We construct a benchmark of 866 instances derived from SV-COMP and evaluate 9 state-of-the-art LLMs across multiple model families. We demonstrate that supervised fine-tuning and Best-of-N sampling yield measurable improvements, and we show that Quokka consistently outperforms prior LLM-based verifiers. Our code and data are publicly available at https://github.com/Anjiang-Wei/Quokka
footnotetext: Correspondence to: anjiang@cs.stanford.edu
@@ 1 Introduction
Loop invariants are conditions that hold before and after each loop iteration, and they are central to deductive program verification. To accelerate program verification, loop invariants must not only be correct but also sufficiently strong to prove the assertions. Generating correct invariants is relatively easy, since any universally true condition qualifies. However, only strong invariants can reduce verification effort and lead to a speedup. For example, in Figure 1, the invariant $x>0$ is correct but not strong enough to prove the final assertion $x\neq 700$, whereas $x\equiv 3\pmod{7}$ is both correct and sufficiently strong.
@@
A series of follow-up works have proposed LLM-based verifiers. Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs. This line of work predominantly adopts complex algorithms to handle LLM-generated invariants, treating LLM output as noisy symbolic material that must be algorithmically repaired or reassembled. For instance, LaM4Inv (Wu et al., 2024a) uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through complex logical operations (conjunctions or disjunctions) to construct valid loop invariants, rather than using them directly. Clause2Inv (Cao et al., 2025) extends this idea further. LOOPY (Kamath et al., 2024) employs a customized Houdini algorithm with iterative refinement to filter candidate invariants. LEMUR (Wu et al., 2024b) develops a recursive and backtracking algorithm for fixing invariants. While these approaches demonstrate the potential of LLM-based verification, their complexity raises a fundamental question: as models become increasingly capable, do we really need such elaborate, harness-heavy designs for building LLM-based verifiers? What if we stop rescuing LLM-generated invariants and instead measure directly whether a candidate invariant helps the solver?
@@
Figure 1: Illustration of Quokka’s evaluation pipeline. The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3). The verification algorithm then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
In this work, we introduce Quokka, an evaluation-oriented framework for LLM-based invariant synthesis that accelerates program verification. Prior LLM-based verification systems treat LLM outputs as imperfect intermediate artifacts requiring substantial symbolic post-processing—predicate filtering, clause combination, Houdini pruning, repair, or backtracking. In contrast, Quokka adopts a simpler, evaluation-centric design: it directly evaluates LLM-generated invariants by checking whether each invariant is valid and whether it helps prove the target assertion. This design eliminates most customized post-processing logic, enables parallel verification queries, and shifts the objective from merely finding correct invariants to finding invariants that can accelerate verification. Despite its simplicity, Quokka achieves state-of-the-art performance compared to all prior LLM-based verifiers.
To support systematic comparison across solvers and LLMs, we construct a dataset of 866 instances derived from the most recent edition of the software verification competition SV-COMP (Beyer and Strejček, 2025). To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date. After evaluating 9 LLMs, we find that Quokka proposes a meaningful and challenging setup for LLMs, where model capabilities can be effectively distinguished and there remains significant room for improvement on challenging cases.
@@
Beyond model evaluation, we investigate techniques to enhance invariant generation capabilities. We construct a training dataset of 3589 instances using a verifier-based filtering approach that ensures data quality. Our experiments demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
@@
  • We introduce a method for LLM-based invariant synthesis that provides sound evaluation and achieves the best performance results compared with prior LLM-based verifiers.
  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 LLMs across multiple model families.
@@ 2 Related Work
A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan and Qadeer, 2002; Lahiri and Bryant, 2007; Hojjat and Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot and Cousot, 1977; Cousot and Halbwachs, 1978; Cousot and Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala and McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich and Bodík, 2018). Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019). Daikon executes the program, observes runtime values, and reports properties that consistently hold over the observed executions. However, such invariants may fail to generalize to all possible executions, thereby compromising soundness. Our approach instead employs a verifier-based algorithm relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
@@
Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014; 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020). More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a; b; c), giving rise to a series of work that explore using LLMs for finding invariants. Pei et al. (2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation. Various techniques have been proposed to couple LLMs with symbolic solvers, including ranking LLM-generated invariants (Chakraborty et al., 2023), the “query-filter-reassemble” strategy of LaM4Inv (Wu et al., 2024a), the back-tracking algorithm in LEMUR (Wu et al., 2024b), and Loopy’s integration of the classic Houdini algorithm (Kamath et al., 2024). On the dataset side, Liu et al. (2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach. In contrast, our work provides a simple and sound evaluation algorithm for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
@@ 3.1 Preliminary
3.2 Verifier-Based Approach
We formalize our verifier-based algorithm for assessing candidate invariants. Let $P$ denote a program. A property is written as $p=\langle\varphi,\ell\rangle$, where $\varphi$ is a state predicate and $\ell$ is a program location. For a finite set $A$ of properties, let $\mathrm{Asm}(P,A)$ be the program obtained from $P$ by inserting assume statements for all elements of $A$. An execution of $\mathrm{Asm}(P,A)$ that reaches a location where an assumption is violated terminates immediately. We write $P\models_{A}p$ to indicate that all executions of $\mathrm{Asm}(P,A)$ satisfy the assertion $p$. The notation $P\models p$ abbreviates $P\models_{\varnothing}p$. Since assumptions restrict behaviors, if $P\not\models_{A}p$ for some $A$, then necessarily $P\not\models p$.
@@ 3.2 Verifier-Based Approach
The verification task specifies a target property $p^{\star}=\langle\varphi^{\star},\ell^{\star}\rangle$. Given $P$ and $p^{\star}$, a large language model proposes a candidate invariant $q=\langle\psi,\ell\rangle$, typically at a loop header. To evaluate the utility of $q$, the algorithm issues two verifier queries:
@@
The query $d_{a}$ verifies whether $q$ is a correct invariant, while the query $d_{b}$ determines whether the target property holds under the assumption that $q$ is true. An illustration of these two verifier queries is provided in Figure 1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure). The outcome of the algorithm is expressed as a judgment
@@
We provide the proof in Appendix A.1. This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct. The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
3.3 Implementation
We describe the implementation of our verifier-based evaluation framework. Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the approach. Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
Syntactic Validation.
Parallel Verifier Queries. For each candidate $q$, the algorithm issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$. These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification. The final outcome is then derived exactly according to the calculus.
3.4 Supervised Fine-Tuning and Best-of-N Sampling
Synthetic Dataset Generation.
We prompt GPT-4o (template in Appendix A.3) with three seed programs to synthesize new C programs containing loops and assertions. Seed programs are randomly drawn from SV-COMP (Beyer and Strejček, 2025), disjoint from our evaluation set to prevent data leakage. Generated programs may fail to compile, contain invalid assertions, or include multiple assertions. We split multi-assertion programs into separate instances, each with a single assertion and all loops. We then run UAutomizer on each program and discard non-compilable instances or those with invalid assertions, yielding 3589 synthetic programs.
Extract Invariants Generated from UAutomizer.
We extract loop invariants from UAutomizer’s output when proving assertions in the synthetic programs. Each invariant includes its program location and predicate. Since programs may contain multiple loops, we pair each program with each of its loop invariants to form our training dataset. An example is shown in Appendix A.5. We perform supervised fine-tuning using LoRA (Hu et al., 2022) on these ground-truth invariants.
Best-of-N Sampling.
Best-of-N sampling generates multiple candidates and selects the most effective one (Ehrlich et al., 2025). We choose the invariant yielding the largest speedup. Since each candidate requires two parallel verifier queries (Section 3.2), Best-of-N evaluates 2N queries concurrently.
4 Experimental Setup
Dataset.
We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard software verification competition, focusing on problems requiring loop invariant synthesis. By filtering instances containing loop keywords (while or for), we obtain 866 instances. To the best of our knowledge, this is the largest dataset used in prior LLM-based verifier evaluations. As shown in Table 1, our programs are significantly longer than prior work. Manual inspection reveals our dataset contains multiple loops, functions, arrays, and pointers (features largely absent from prior datasets), making it a more challenging benchmark that better distinguishes solver performance. For training data, we construct 3589 instances with synthetic data generation filtered by the solvers, as described in Section 3.4.
Dataset Avg. #Lines #Instances Multi-Loops Multi-Functions Arrays/Pointers
LaM4Inv (Wu et al., 2024a) 18 316
Clause2Inv (Cao et al., 2025) 18 366
LEMUR (Wu et al., 2024b) 20 180
Loopy (Kamath et al., 2024) 25 638
Quokka (Ours) 52 866
Table 1: Dataset statistics.
Metrics.
We evaluate LLMs on two dimensions: (1) correctness of generated invariants and (2) performance improvements. Correctness is verified by the solver with a timeout of $1.2\times$ the original solving time, providing slack to reduce inconclusive outcomes. We report the number of additionally solved instances (unsolvable by the base solver), total solved instances, and average solving time under varying timeout budgets. All measurements represent end-to-end execution time, including LLM inference overhead.
Choice of Solvers.
We use UAutomizer (Schüssele et al., 2024), the state-of-the-art non-LLM-based solver, as our base solver. Our methodology requires that the solver can verify programs without externally-provided invariants, enabling speedup measurement relative to a functional baseline. We select UAutomizer over alternatives like Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018) because it represents the current state-of-the-art (outperforming ESBMC on SV-COMP benchmarks) and, unlike Frama-C, has built-in invariant synthesis capabilities for autonomous verification. To demonstrate generality, we also present ESBMC experiments in Appendix A.8.
Additional details such as models and hardware setup are provided in Appendix A.2.
5 Results
5.1 Results of LLMs
Model # Correct Invariants 60s 500s
$\Delta$ # Solved $\overline{T}$ (s) $\Delta$ # Solved $\overline{T}$ (s)
Llama-3.1-8B 342 5 321 41.9 0 342 309.9
Qwen2.5-7B 419 5 402 37.6 0 417 267.5
claude-opus-4.1 487 13 466 34.2 1 487 229.4
Qwen2.5-72B 500 11 474 33.8 0 500 221.7
o3 550 17 534 35.9 1 549 198.3
claude-sonnet-4 620 16 591 27.8 1 620 156.0
claude-opus-4.5 689 17 659 24.2 1 689 118.8
gpt-5.1 694 15 661 23.3 1 694 113.8
gpt-5.2 710 21 681 22.2 1 710 105.1
Table 2: Performance comparison of LLMs on invariant generation: number of correct invariants, additional instances solved beyond UAutomizer baseline ($\Delta$), total solved instances, and average solving time ($\overline{T}$) under 60s and 500s timeouts.
Model # Correct Invariants 60s 500s
# $\Delta$ # Solved $\overline{T}$ # $\Delta$ # Solved $\overline{T}$
Qwen2.5-7B (base) 419 5 402 37.6 0 417 267.5
Qwen2.5-7B (finetuned) 431 7 414 36.8 0 429 260.8
Table 2 presents the performance of different LLMs when using Quokka. Overall, the table demonstrates substantial variation in both correctness and performance across different models, with larger and more recent models generally achieving better results.
Quokka proposes a meaningful and challenging problem setup for LLMs.
The results in Table 2 reveal an interesting performance pattern across timeout budgets. Under the 60s timeout, LLMs solve much more instances beyond the UAutomizer baseline (ranging from 5 additional instances for smaller models to 21 for gpt-5.2), demonstrating their ability to accelerate verification on moderately difficult problems. However, under the 500s timeout, the incremental benefit diminishes significantly, indicating that LLMs struggle to provide meaningful invariants for the hardest problems that require more solving time. This pattern suggests that Quokka establishes a well-calibrated benchmark: LLMs already provide meaningful acceleration on a portion of instances, yet significant room remains for improvement on challenging cases. The substantial variation in correct invariants (342 for Llama-3.1-8B to 710 for gpt-5.2) and average solving times (41.9s to 22.2s under 60s timeout) further demonstrates that the benchmark effectively distinguishes model capabilities. While alternative verification algorithms might yield different absolute performance numbers, the superiority and simplicity of Quokka over other LLM-based verifiers (discussed in Section 5.4) establishes it as an appropriate framework for evaluating LLMs’ invariant generation capabilities.
We provide further analysis of model inference time in Appendix A.7 and evaluate alternative prompting strategies (few-shot and chain-of-thought prompting) in Appendix A.9.
5.2 Failure Mode Analysis
To understand why many invariants fail to accelerate verification in Quokka, we analyze the failure modes of the top three models, with detailed results provided in Appendix A.11. The analysis shows that although some failures are caused by incorrect invariants, the majority are due to timeouts.
We randomly sampled 50 timeout cases and manually analyzed the invariants. Our analysis reveals several key patterns. First, only 7 cases (14%) involved genuinely incorrect invariants, indicating reasonable LLM correctness. Second, 19 cases (38%) generated invariants identical to the original assertion, providing no decomposition benefit, with 4 additional cases nearly identical. Third, 9 cases (18%) generated trivial invariants (e.g., simple bounds) that failed to capture essential loop behavior. Fourth, only 5 cases (10%) produced strong, non-trivial invariants that still timed out, suggesting inherently difficult verification queries. Finally, 6 cases are incorrect original assertions where the LLM generated correct invariants. We provide case studies in Appendix A.12.
5.3 Results of Fine-Tuning and Best-of-N Sampling
Figure 2: Effect of Best-of-N sampling on the number of $\Delta$ instances solved over Uautomizer under different timeout settings.
Modest Improvement from Supervised Fine-Tuning.
We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs. As shown in Table 3, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, additionally solved instances increase from 5 to 7 under 60s timeout. These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
Benefits of Best-of-N Sampling.
We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time. As shown in Figure 2, increasing N from 1 to 4 generally improves performance. The benefits continue to grow up to N=8, which achieves the best results. Beyond N=8, performance plateaus as resource contention and memory constraints begin to offset the advantages of additional sampling. These results demonstrate that Best-of-N sampling is an effective strategy for improving verification performance, with optimal gains observed at N=8 in our experimental setup.
5.4 Comparison with Prior LLM-based Verifiers
Method 30s 60s 250s 350s 500s
# $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s) # $\Delta$ $\overline{T}$(s)
LaM4Inv 1 30.0 1 59.9 1 249.3 0 348.9 0 498.4
Clause2Inv-original 14 25.8 9 49.9 4 199.2 2 277.4 0 394.4
Clause2Inv-gpt-5.2 20 25.1 13 48.2 6 190.7 3 265.2 1 377.0
Loopy 0 30.0 0 60.0 5 240.0 3 299.5 2 360.4
LEMUR-original 7 28.6 7 50.4 3 165.3 1 222.3 1 306.8
LEMUR-gpt-5.2 15 27.9 13 47.9 7 153.9 4 205.9 2 282.8
Quokka-gpt-5.2 (best-of-8) 41 17.5 25 26.6 7 78.5 4 104.8 3 143.8
Quokka-gpt-5.2 (2-shot) 33 14.3 20 19.4 6 45.2 2 57.6 2 76.0
Table 4: Extra solved instances over UAutomizer and average solving time at different timeout thresholds.
Table 4 compares Quokka against prior LLM-based verifiers, reporting additional instances solved beyond UAutomizer and average solving time under different timeout budgets. Figure 3 shows cumulative solved instances over time, demonstrating Quokka’s superior performance across all timeout thresholds.
Method 30s 60s 250s 350s 500s
# Solved $\overline{T}$(s) # Solved $\overline{T}$(s) # Solved $\overline{T}$(s) # Solved $\overline{T}$(s) # Solved $\overline{T}$(s)
LEMUR-gpt-5.2 4 27.9 9 34.0 18 75.9 19 88.1 20 106.6
Quokka-gpt-5.2 19 9.7 20 11.6 24 28.8 24 28.8 24 28.8
Table 5: Performance comparison on the benchmark used in LEMUR. Since UAutomizer solves 0 instances within ten minutes on this dataset, # Solved equals # $\Delta$ (additional instances solved beyond the baseline). Results shown for different timeout thresholds.
Quokka Outperforms Prior LLM-based Verifiers.
Quokka with gpt-5.2 and best-of-8 sampling achieves state-of-the-art results across all speedup metrics, while Quokka with gpt-5.2 and 2-shot prompting also achieves leading performance under most timeout thresholds. Among prior work, LEMUR (Wu et al., 2024b) and Clause2Inv (Cao et al., 2025) represent the strongest baselines. LEMUR supports both UAutomizer and ESBMC. We evaluate LEMUR with UAutomizer, which is the state-of-the-art solver and matches the base solver used in our method. For fair comparison, we upgrade both LEMUR and Clause2Inv to gpt-5.2 (from gpt-4 and gpt-4o-mini, respectively), creating LEMUR-gpt-5.2 and Clause2Inv-gpt-5.2. Quokka-gpt-5.2 outperforms both upgraded baselines. LaM4Inv (Wu et al., 2024a) [^1] LaM4Inv’s performance differs from its original paper due to dataset differences and its requirement for manual preprocessing (confirmed by the authors), for which no automated script exists. and LOOPY (Kamath et al., 2024) show minimal gains over UAutomizer.
To further quantify the performance gap at larger timeout, we evaluate on LEMUR’s benchmark (Table 5)—challenging instances where UAutomizer fails within 10 minutes. Quokka consistently outperforms LEMUR-gpt-5.2 across all timeouts, solving substantially more instances with lower average solving times.
Analysis.
The baselines differ in both base solver choices and algorithm design. We analyze base verifier differences in Appendix A.10. Regarding algorithm design, Quokka adopts a distinctly “evaluation-oriented” approach: rather than implementing a handcrafted invariant search algorithm, it provides a clean interface between an LLM and a verifier. This represents a conceptual shift from prior work, which introduces complex algorithms to transform imperfect LLM outputs into usable invariants. Prior approaches treat LLM output as noisy symbolic material requiring algorithmic repair, whereas Quokka treats it as a direct proof candidate to be validated by the verifier. For instance, LaM4Inv requests complete invariants but then filters predicates with model checker and reassembles them, indicating distrust in the raw output. Clause2Inv takes this further. However, as models become increasingly capable, simpler evaluation-centric designs become preferable. We additionally provide a detailed comparison between Quokka and LEMUR in Appendix A.10.
6 Conclusion
This work introduced Quokka, an evaluation-oriented framework for LLM-based invariant synthesis that accelerates program verification. Unlike prior work that treats LLM outputs as noisy symbolic material requiring substantial post-processing, Quokka adopts a simpler and evaluation-centric design that directly validates whether each LLM-generated invariant helps prove the target assertion. We constructed a benchmark of 866 instances derived from SV-COMP and evaluated 9 state-of-the-art LLMs across multiple model families. Our results demonstrate that Quokka consistently outperforms prior LLM-based verifiers. We further showed that supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification. Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
Appendix A Appendix
A.1 Soundness Proof
Theorem (Decision Soundness).